OpenStudy (yumyum247):

how do i write this????? common log of the square root of 9/25.

OpenStudy (yumyum247):

@jim_thompson5910

OpenStudy (zzr0ck3r):

\(\log(\sqrt{\dfrac{9}{25}})=\log(\dfrac{\sqrt{9}}{\sqrt{25}})=\log(\dfrac{3}{5})\)

OpenStudy (yumyum247):

OH OK THANKS ZZROCK ☺😁😁

OpenStudy (zzr0ck3r):

or \(\log(\sqrt{\dfrac{9}{25}})=\log((\dfrac{9}{25})^\frac{1}{2})=\dfrac{1}{2}\log(\frac{9}{25})\)

OpenStudy (eli-ite):

um i don't know i need a calculator